# X-Dreamer: Creating High-quality 3D Content by Bridging the Domain Gap Between Text-to-2D and Text-to-3D Generation

### v2

@@ 3.1 Architecture
where $\mathcal{P}(\cdot)$ represents the positional encoding using a hash-grid technique [20]. Subsequently, each pixel of the rendered image can be computed as follows:
$V(p,\omega)=\int_{\Omega}L_{i}(p,\omega_{i})f(p,\omega_{i},\omega)(\omega_{i}\cdot n_{p})\mathrm{d}\omega_{i},$ (6)
@@
where $V(p,\omega)$ denotes the rendered pixel value from the direction $\omega$ for the surface point $p$. $\Omega$ denotes a hemisphere defined by the set of incident directions $\omega_{i}$ satisfying the condition $\omega_{i}\cdot n_{p}\geq 0$, where $\omega_{i}$ denotes the incident direction, and $n_{p}$ represents the surface normal at point $p$. $L_{i}(\cdot)$ corresponds to the incident light from an off-the-shelf environment map, and $f(\cdot)$ is the Bidirectional Reflectance Distribution Function (BRDF) related to the material properties (i.e., $\boldsymbol{k}_{d},\boldsymbol{k}_{n},\boldsymbol{k}_{rm}$). By aggregating all rendered pixel colors, we obtain a rendered image $\boldsymbol{x}=\{V(p,\omega)\}$. Similar to the geometry modeling stage, we feed the rendered image $\boldsymbol{x}$ into SD. The optimization objective remains the same as Eq. 3 and Eq. 4, where the rendered normal map $\boldsymbol{n}$ and the parameters of DMTet $\Phi_{det}$ are replaced with the rendered image $\boldsymbol{x}$ and the parameters of the material encoder $\Phi_{mat}$, respectively.
3.2 Camera-Guided Low-Rank Adaptation
@@
The domain gap between text-to-2D and text-to-3D generation presents a significant challenge, as discussed in Sec. 1. It has been observed that directly utilizing pretrained SD for text-to-3D generation can result in certain issues, such as the Janus Problem [12, 1]. To address these issues, we propose Camera-Guided Low-Rank Adaptation (CG-LoRA) as a solution to bridge the domain gap. As depicted in Fig. 20, we leverage camera parameters and direction-aware text to guide the generation of parameters in CG-LoRA, enabling X-Dreamer to effectively incorporate camera perspective and direction information.
@@
$\boldsymbol{t}$ $=\mathcal{E}_{txt}(T),$ (7)
$\boldsymbol{c}$ $=\mathcal{E}_{pos}(C),$ (8)
@@
$\boldsymbol{A}_{txt}$ $=\text{Reshape}(\boldsymbol{t}\boldsymbol{W}_{txt}),$ (9)
$\boldsymbol{A}_{cam}$ $=\text{Reshape}(\boldsymbol{c}\boldsymbol{W}_{cam}),$ (10)
@@
$\boldsymbol{y}=\boldsymbol{x}\boldsymbol{W}+[\boldsymbol{x}\boldsymbol{A}_{txt};\boldsymbol{x}\boldsymbol{A}_{cam}]\boldsymbol{B},$ (11)
@@
Figure 4: Text-to-3D generation results from an ellipsoid.
Figure 5: Text-to-3D generation results from coarse-grained guided meshes.
Figure 6: Comparison with State-of-the-Art (SOTA) methods. Our method yields results that exhibit enhanced fidelity and more details.
3.3 Attention-Mask Alignment Loss
@@
$\bar{\boldsymbol{a}}=\mathrm{Softmax}(\frac{\boldsymbol{Q}\boldsymbol{K}^{\top}}{\sqrt{d}}),$ (12)
@@
where $H$ denotes the number of attention heads in SD, and $\bar{\boldsymbol{a}}\in\mathbb{R}^{H\times h\times w}$ represents the attention map. Subsequently, we proceed to compute the overall attention map $\hat{\boldsymbol{a}}\in\mathbb{R}^{h\times w}$ by averaging the attention values of $\bar{\boldsymbol{a}}$ across all attention heads. Since the attention map values are normalized using the softmax function, the activation values in the attention map may become very small when the image feature resolution is high. However, considering that each element in the rendered mask has a binary value of either 0 or 1, directly aligning the attention map with the rendered mask is not optimal. To address this, we propose a normalization technique that maps the values in the attention map from 0 to 1. This normalization process is formulated as follows:
$\boldsymbol{a}=\frac{\hat{\boldsymbol{a}}-\min(\hat{\boldsymbol{a}})}{\max(\hat{\boldsymbol{a}})-\min(\hat{\boldsymbol{a}})+\nu},$ (13)
@@
where $\nu$ represents a small constant value (e.g., $-61e\textrm{-}$ 6) that prevents division by zero in the denominator. Finally, we align the attention maps of all attention layers with the rendered mask of the 3D object using the AMA loss. The formulation of this alignment is presented in Eq. 4.
4 Experiments
4.1 Implementation Details.
@@ 4 Experiments
We conduct the experiments using four Nvidia RTX 3090 GPUs and the PyTorch library [21]. To calculate the SDS loss, we utilize the Stable Diffusion implemented by HuggingFace Diffusers [34]. For the DMTet $\Phi_{dmt}$ and material encoder $\Phi_{mat}$, we implement them as a two-layer MLP and a single-layer MLP, respectively, with a hidden dimension of 32. The values of $d_{cam}$, $d_{txt}$, $r$, $r^{\prime}$, the batch size, the SDS loss weight, the AMA loss weight, and the aspect ratio of the perspective projection plane are set to 1024, 1024, 4, 4, 4, 1, 0.1, and 1 respectively. We optimize X-Dreamer for 2000 iterations for geometry learning and 1000 iterations for appearance learning. For each iteration, $\phi_{pit}$, $\phi_{yaw}$, and $\theta_{fov}$ are randomly sampled from $(-15^{\circ},45^{\circ})$, $(-180^{\circ},180^{\circ})$, and $(25^{\circ},45^{\circ})$, respectively.
4.2 Results of X-Dreamer
@@
Text-to-3D generation from an ellipsoid. We present representative results of X-Dreamer for text-to-3D generation, utilizing an ellipsoid as the initial geometry, as shown in Fig. 4. The results demonstrate the ability of X-Dreamer to generate high-quality and photo-realistic outputs that accurately correspond to the input text prompts.
@@
Text-to-3D generation from coarse-grained meshes. While there is a wide availability of coarse-grained meshes for download from the internet, directly utilizing these meshes for 3D content creation often results in poor performance due to the lack of geometric details. However, when compared to a 3D ellipsoid, these meshes may provide better 3D shape prior information for X-Dreamer. Hence, instead of using ellipsoids, we can initialize DMTet with coarse-grained guided meshes as well. As shown in Fig. 5, X-Dreamer can generate 3D assets with precise geometric details based on the given text, even when the provided coarse-grained mesh lacks details. For instance, in the last column of Fig. 5, X-Dreamer accurately transforms the geometry from a cow to a corgi based on the text prompt “A corgi, highly detailed.” Therefore, X-Dreamer is also an exceptionally powerful tool for editing coarse-grained mesh geometry using textual inputs.
Figure 7: Ablation studies of the proposed X-Dreamer.
Figure 8: Visualization of Attention Map, Rendered Mask, and Rendered Image with and without AMA Loss. For clarity, we only visualize the attention map of the first attention layer in SD.
@@
Qualitative Comparison. To assess the effectiveness of X-Dreamer, we compare it with four SOTA methods: DreamFusion [22], Magic3D [13], Fantasia3D [4], and ProlificDreamer [36], as depicted in Fig. 6. When compared to the SDS-based methods [22, 13, 36], X-Dreamer outperforms them in generating superior-quality and realistic 3D assets. In addition, when compared to the VSD-based method [36], X-Dreamer produces 3D content with comparable or even better visual effects, while requiring significantly less optimization time. Specifically, the geometry and appearance learning process of X-Dreamer requires only approximately 27 minutes, whereas ProlificDreamer exceeds 8 hours.
4.3 Ablation Study
@@
Ablation on the proposed modules. To gain insights into the abilities of CG-LoRA and AMA loss, we perform ablation studies wherein each module is incorporated individually to assess its impact. As depicted in Fig. 7, the ablation results demonstrate a notable decline in the geometry and appearance quality of the generated 3D objects when CG-LoRA is excluded from X-Dreamer. For instance, as shown in the second row of Fig. 7, the generated Batman lacks an ear on the top of its head in the absence of CG-LoRA. This observation highlights the crucial role of CG-LoRA in injecting camera-relevant information into the model, thereby enhancing the 3D consistency. Furthermore, the omission of AMA loss from X-Dreamer also has a deleterious effect on the geometry and appearance fidelity of the generated 3D assets. Specifically, as illustrated in the first row of Fig. 7, X-Dreamer successfully generates a photorealistic texture for the rocket, whereas the texture quality noticeably deteriorates in the absence of AMA loss. This disparity can be attributed to AMA loss, which directs the focus of the model towards foreground object generation, ensuring the realistic representation of both geometry and appearance of foreground objects. These ablation studies provide valuable insights into the individual contributions of CG-LoRA and AMA loss in enhancing the geometry, appearance, and overall quality of the generated 3D objects.
@@
Attention map comparisons w/ and w/o AMA loss. AMA loss is introduced with the aim of guiding attention during the denoising process towards the foreground object. This objective is achieved by aligning the attention map of SD with the rendered mask of the 3D object. To evaluate the effectiveness of AMA loss in accomplishing this goal, we visualize the attention maps of SD with and without AMA loss at both the geometry learning and appearance learning stages. As depicted in Fig. 8, it can be observed that incorporating AMA loss not only results in improved geometry and appearance of the generated 3D asset, but also concentrates the attention of SD specifically on the foreground object area. The visualizations confirm the efficacy of AMA loss in directing the attention of SD, resulting in improved quality and foreground object focus during geometry and appearance learning stages.
5 Conclusion

### v3

@@ 3.1.2. Appearance Learning
where $\mathcal{P}(\cdot)$ represents the positional encoding using a hash-grid technique (Müller et al., 2022). Subsequently, each pixel of the rendered image can be computed as follows:
(6) $V(p,\omega)=\int_{\Omega}L_{i}(p,\omega_{i})f(p,\omega_{i},\omega)(\omega_{i}\cdot n_{p})\mathrm{d}\omega_{i},$
@@
where $V(p,\omega)$ denotes the rendered pixel value from the direction $\omega$ for the surface point $p$. $\Omega$ denotes a hemisphere defined by the set of incident directions $\omega_{i}$ satisfying the condition $\omega_{i}\cdot n_{p}\geq 0$, where $\omega_{i}$ denotes the incident direction, and $n_{p}$ represents the surface normal at point $p$. $L_{i}(\cdot)$ corresponds to the incident light from an off-the-shelf environment map, and $f(\cdot)$ is the Bidirectional Reflectance Distribution Function (BRDF) related to the material properties (i.e., $\boldsymbol{k}_{d},\boldsymbol{k}_{n},\boldsymbol{k}_{rm}$). By aggregating all rendered pixel colors, we obtain a rendered image $\boldsymbol{x}=\{V(p,\omega)\}$. Similar to the geometry modeling stage, we feed the rendered image $\boldsymbol{x}$ into SD. The optimization objective remains the same as Equ. 3 and Equ. 4, where the rendered normal map $\boldsymbol{n}$ and the parameters of DMTet $\Phi_{dmt}$ are replaced with the rendered image $\boldsymbol{x}$ and the parameters of the material encoder $\Phi_{mat}$, respectively.
3.2. Camera-Guided Low-Rank Adaptation
@@
The domain gap between text-to-2D and text-to-3D generation presents a significant challenge, as discussed in Sec. 1, which leads to unsatisfied generated results. To address these issues, we propose Camera-Guided Low-Rank Adaptation (CG-LoRA) as a solution to bridge the domain gap. As depicted in Fig. 4, we leverage camera parameters and direction-aware text to guide the generation of parameters in CG-LoRA, enabling X-Dreamer to effectively incorporate camera perspective and direction information.
@@
(7) $\boldsymbol{t}$ $=\mathcal{E}_{txt}(T),$
(8) $\boldsymbol{c}$ $=\mathcal{E}_{pos}(C),$
@@
(9) $\boldsymbol{A}_{txt}$ $=\text{Reshape}(\boldsymbol{t}\boldsymbol{W}_{txt}),$
(10) $\boldsymbol{A}_{cam}$ $=\text{Reshape}(\boldsymbol{c}\boldsymbol{W}_{cam}),$
@@
(11) $\boldsymbol{y}=\boldsymbol{x}\boldsymbol{W}+[\boldsymbol{x}\boldsymbol{A}_{txt};\boldsymbol{x}\boldsymbol{A}_{cam}]\boldsymbol{B},$
@@
3.3. Attention-Mask Alignment Loss
@@
(12) $\bar{\boldsymbol{a}}=\mathrm{Softmax}(\frac{\boldsymbol{Q}\boldsymbol{K}^{\top}}{\sqrt{d}}),$
@@
where $H$ denotes the number of attention heads in SD. $h$, $w$, and $d$ represent the height, width, and channel dimensions of the image features, respectively. $\bar{\boldsymbol{a}}\in\mathbb{R}^{H\times h\times w}$ represents the attention map. Subsequently, we proceed to compute the overall attention map $\hat{\boldsymbol{a}}\in\mathbb{R}^{h\times w}$ by averaging the attention values of $\bar{\boldsymbol{a}}$ across all attention heads. Since the attention map values are normalized using the softmax function, the activation values in the attention map may become very small when the image feature resolution is high. However, considering that each element in the rendered mask has a binary value of either 0 or 1, directly aligning the attention map with the rendered mask is not optimal. To address this, we propose a normalization technique that maps the values in the attention map from 0 to 1. This normalization process is formulated as follows:
(13) $\boldsymbol{a}=\frac{\hat{\boldsymbol{a}}-\min(\hat{\boldsymbol{a}})}{\max(\hat{\boldsymbol{a}})-\min(\hat{\boldsymbol{a}})+\nu},$
@@
where $\nu$ represents a small constant value (e.g., $-61e\textrm{-}$ 6) that prevents division by zero in the denominator. Finally, we align the attention maps of all attention layers with the rendered mask of the 3D object using the AMA loss. The formulation of this alignment is presented in Equ. 4.
4. Experiments
4.1. Implementation Details.
@@ 4. Experiments
We conduct the experiments using four Nvidia RTX 3090 GPUs and the PyTorch library (Paszke et al., 2019). To calculate the SDS loss, we utilize the Stable Diffusion implemented by HuggingFace Diffusers (von Platen et al., 2022). For the DMTet $\Phi_{dmt}$ and material encoder $\Phi_{mat}$, we implement them as a two-layer MLP and a single-layer MLP, respectively, with a hidden dimension of 32. The values of $d_{cam}$, $d_{txt}$, $r$, $r^{\prime}$, the batch size, the SDS loss weight, the AMA loss weight, and the aspect ratio of the perspective projection plane are set to 1024, 1024, 4, 4, 4, 1, 0.1, and 1 respectively. AMA loss is used after half of the training iterations, where CG-LoRA has a certain degree of perspective alignment ability. We optimize X-Dreamer for 2000 iterations for geometry learning and 1000 iterations for appearance learning. For each iteration, $\phi_{pit}$, $\phi_{yaw}$, and $\theta_{fov}$ are randomly sampled from $(-15^{\circ},45^{\circ})$, $(-180^{\circ},180^{\circ})$, and $(25^{\circ},45^{\circ})$, respectively.
Figure 5. Text-to-3D generation results from an ellipsoid.
Figure 6. Text-to-3D generation results from coarse-grained guided meshes.
4.2. Results of X-Dreamer
4.2.1. Text-to-3D generation from an ellipsoid
@@
We present representative results of X-Dreamer for text-to-3D generation, utilizing an ellipsoid as the initial geometry, as shown in Fig. 5. The results demonstrate the ability of X-Dreamer to generate high-quality and photo-realistic outputs that accurately correspond to the input text prompts.
4.2.2. Text-to-3D generation from coarse-grained meshes
@@
While there is a wide availability of coarse-grained meshes for download from the internet, directly utilizing these meshes for 3D content creation often results in poor performance due to the lack of geometric details. However, when compared to a 3D ellipsoid, these meshes may provide better 3D shape prior information for X-Dreamer. Hence, instead of using ellipsoids, we can initialize DMTet with coarse-grained guided meshes as well. As shown in Fig. 6, X-Dreamer can generate 3D assets with precise geometric details based on the given text, even when the provided coarse-grained mesh lacks details. For instance, in the last column of Fig. 6, X-Dreamer accurately transforms the geometry from a cow to a corgi based on the text prompt “A corgi, highly detailed.” Therefore, X-Dreamer is also an exceptionally powerful tool for editing coarse-grained mesh geometry using textual inputs.
Table 1. Quantitative comparison of SOTA Methods: The top-performing and second-best results are highlighted in bolded and underlined, respectively.
Method User Study CLIP Score OpenCLIP Score
Geo. Qua. App. Qua. ViT-B/32 ViT-B/16 ViT-L/14 ViT-B/32 ViT-B/16 ViT-L/14
Dreamfusion (Poole et al., 2022) 1.2 0.6 30.5 31.1 26.3 31.9 27.9 30.0
Magic3d (Lin et al., 2023) 2.1 1.1 28.4 28.6 24.6 29.1 25.9 28.3
Fantasia3d (Chen et al., 2023a) 5.4 6.3 30.4 30.1 24.8 30.3 27.2 29.7
[…]
X-Dreamer 78.4 75.8 32.0 32.2 27.2 34.1 29.6 32.2
Figure 7. Comparison with SOTA methods. Our method yields results that exhibit enhanced fidelity and more details.
Figure 8. More comparison with State-of-the-Art (SOTA) methods.
4.2.3. Quantitative Comparison
We conducted a meticulously organized user study involving 50 individuals to evaluate the performance of four state-of-the-art (SOTA) methods by assessing 50 generated results, i.e., DreamFusion (Poole et al., 2022), Magic3D (Lin et al., 2023), Fantasia3D (Chen et al., 2023a), and ProlificDreamer (Wang et al., 2023c). Participants were asked to compare and select the best result based on two criteria: Geometry Quality (Geo. Qua.) and Appearance Quality (App. Qua.). The results, as presented in Tab. 1, indicate that our approach garnered a preference from 75.0% of the participants. This overwhelming preference underscores our method’s exceptional quality and superiority, establishing it as a pioneering solution in the field. The primary objective of these user studies was to evaluate the geometric and visual quality of the generated results, without specifically assessing their alignment with the text. To provide a more comprehensive evaluation, we further calculated the CLIP score and OpenCLIP score for both the generated results and the corresponding text prompts. As depicted in Tab. 1, our method outperformed other approaches in terms of both the CLIP score and OpenCLIP score. These findings demonstrate that the results generated by our method exhibit superior alignment with the text, further strengthening the efficacy of our approach.
4.2.4. Qualitative Comparison
@@
To assess the effectiveness of X-Dreamer, we compare it with four SOTA methods. Since the codes of some methods (Poole et al., 2022; Lin et al., 2023; Wang et al., 2023c) have not been publicly available, we present the results obtained by implementing these methods in threestudio (Guo et al., 2023). The results are depicted in Fig. 7. In Fig. 8, we present a comprehensive comparison of our methods with four baselines, using the images provided in their original papers. When compared to the SDS-based methods (Poole et al., 2022; Lin et al., 2023; Chen et al., 2023a), X-Dreamer outperforms them in generating superior-quality and realistic 3D assets. In addition, when compared to the VSD-based method (Wang et al., 2023c), X-Dreamer produces 3D content with comparable or even better visual effects, while requiring significantly less optimization time. Specifically, the geometry and appearance learning process of X-Dreamer requires only approximately 27 minutes, whereas ProlificDreamer exceeds 8 hours.
Figure 9. Camera Control Comparison with Zero-1-to-3 (Liu et al., 2023).
4.2.5. Camera Control Comparison
During training, CG-LoRA aims to control the camera parameters. To investigate the effectiveness of this control, we visualize the image generated by Stable Diffusion (SD) with a pretrained CG-LoRA. As shown in Fig. 9, given a specific camera parameter, SD with CG-LoRA can generate images from that perspective. To make a comparative analysis, we also compare our results with Zero-1-to-3 (Liu et al., 2023), a method capable of generating images with desired viewing angles based on relative camera angles and reference images. Using the front view produced by our approach as a reference image, we leverage Zero-1-to-3 to generate images with varying camera parameters. Our observations indicate that our method can generate images that exhibit a general alignment with the angles produced by Zero-1-to-3.
4.3. Ablation Study
Figure 10. Ablation studies of the proposed X-Dreamer.
Figure 11. Visualization of Attention Map, Rendered Mask, and Rendered Image with and without AMA Loss. For clarity, we only visualize the attention map of the first attention layer in SD.
4.3.1. Ablation on the proposed modules
@@
To gain insights into the abilities of CG-LoRA and AMA loss, we perform ablation studies wherein each module is incorporated individually to assess its impact. As depicted in Fig. 10, the ablation results demonstrate a notable decline in the geometry and appearance quality of the generated 3D objects when CG-LoRA is excluded from X-Dreamer. For instance, as shown in the second row of Fig. 10, the generated Batman lacks an ear on the top of its head in the absence of CG-LoRA. This observation highlights the crucial role of CG-LoRA in injecting camera-relevant information into the model, thereby enhancing the 3D consistency. Furthermore, the omission of AMA loss from X-Dreamer also has a deleterious effect on the geometry and appearance fidelity of the generated 3D assets. Specifically, as illustrated in the first row of Fig. 10, X-Dreamer successfully generates a photorealistic texture for the rocket, whereas the texture quality noticeably deteriorates in the absence of AMA loss. This disparity can be attributed to AMA loss, which directs the focus of the model towards foreground object generation, ensuring the realistic representation of both geometry and appearance of foreground objects. These ablation studies provide valuable insights into the individual contributions of CG-LoRA and AMA loss in enhancing the geometry, appearance, and overall quality of the generated 3D objects.
4.3.2. Attention map comparisons w/ and w/o AMA loss
@@
AMA loss is introduced with the aim of guiding attention during the denoising process towards the foreground object. This objective is achieved by aligning the attention map of SD with the rendered mask of the 3D object. To evaluate the effectiveness of AMA loss in accomplishing this goal, we visualize the attention maps of SD with and without AMA loss at both the geometry learning and appearance learning stages. As depicted in Fig. 11, it can be observed that incorporating AMA loss not only results in improved geometry and appearance of the generated 3D asset, but also concentrates the attention of SD specifically on the foreground object area. The visualizations confirm the efficacy of AMA loss in directing the attention of SD, resulting in improved quality and foreground object focus during geometry and appearance learning stages.
4.3.3. Ablation on the parameter generation of CG-LoRA
Figure 12. Ablation Study of CG-LoRA.
In our implementation, the parameters of CG-LoRA are dynamically generated based on two key factors related to camera information: camera parameters and direction-aware text. To thoroughly investigate the impact of these camera information terms on X-Dreamer, we conducted ablation experiments, dynamically generating CG-LoRA parameters based solely on one of these terms. The results, as illustrated in Fig. 12, clearly demonstrate that when using only one type of camera information term, the geometry and appearance quality of the generated results are significantly diminished compared to those produced by X-Dreamer. For example, as shown in the third row of Fig. 12, when utilizing the CG-LoRA solely related to direction-aware text, the geometry quality of the generated cookies appears poor. Similarly, as depicted in the last row of Fig. 12, employing the CG-LoRA solely based on camera parameters results in diminished geometry and appearance quality for the pumpkin. These findings highlight the importance of both camera parameters and direction-aware text in generating high-quality 3D assets. The successful generation of parameters for CG-LoRA relies on the synergy between these two camera information terms, emphasizing their crucial role in achieving superior geometry and appearance in the generated assets.
4.4. Attention Map of Different Layers
Figure 13. Attention maps of different layers of SD for the geometry learning stage.
Figure 14. Attention maps of different layers of SD for the appearance learning stage.
Figure 15. Rendered images of generated 3D assets under different environment maps.
Figure 16. Bad cases of X-Dreamer.
In this section, we present visualizations of the attention maps from various attention layers during the geometry learning stage and the appearance learning stage. As illustrated in Fig. 13 and Fig. 14, the attention maps generated by different attention layers consistently exhibit the ability to accurately locate the shape of the foreground object. This demonstrates the robustness and effectiveness of the attention mechanism in capturing the essential features of the foreground object. Furthermore, we observe that attention maps from different layers seem to exhibit variations in focus, with some emphasizing the foreground, some highlighting the edges, and some concentrating on the interior of the object. This divergence in focus can be attributed to the distinct roles played by different layers in the stable diffusion training process. Each layer contributes its unique perspective and emphasis, leading to variations in attention focus. Nevertheless, despite these variations, all attention maps generated by X-Dreamer successfully locate the foreground objects with precision. This outcome provides strong evidence that our proposed module effectively guides the model’s attention towards the foreground objects. By incorporating the attention mechanism into the training process, X-Dreamer ensures that the model focuses on the most relevant regions, resulting in accurate and reliable detection of the foreground objects throughout the geometry and appearance learning stages.
4.5. Rendered Images under Different Environment Maps
Our method, while employing a fixed environment for training the model, offers significant advantages and flexibility in the rendering of 3D assets generated by X-Dreamer. It is worth noting that these assets are not limited to a single environment; instead, they possess the capability to be rendered in diverse environments, expanding their applicability and adaptability. The visual evidence presented in Fig. 15 reinforces our claim. By utilizing different environment maps, we can readily observe the striking variations in the appearance of the rendered images of these 3D assets. This compelling demonstration showcases the inherent versatility and potential of X-Dreamer’s generated assets to seamlessly integrate into various rendering engines. The ability to effectively address the challenges posed by different lighting conditions is a crucial aspect of any rendering engine. By leveraging the capabilities of X-Dreamer, we unlock the possibility of using its 3D assets in a multitude of rendering engines. This versatility ensures that the assets can adapt and thrive in diverse lighting scenarios, enabling them to meet the demands of real-world applications in fields such as architecture, virtual reality, and computer graphics.
4.6. Limitation
One limitation of the proposed X-Dreamer is its inability to generate multiple objects simultaneously. When the input text prompt contains two distinct objects, X-Dreamer may produce a mixed object that combines properties of both objects. A clearer understanding of this limitation can be gained by referring to Fig. 16. In the first row of Fig. 16, our intention is for X-Dreamer to generate “A sliced loaf of fresh bread and a cabbage.” However, the resulting output consists of a bread with a few cabbage leaves, indicating the model’s difficulty in generating separate objects accurately. Similarly, in the second row of Fig. 16, we aim to generate “A blue and white porcelain vase and an apple.” However, the output depicts a blue and white porcelain vase with a shape resembling an apple. Nevertheless, it is important to note that this limitation does not significantly undermine the value of our work. In applications such as games and movies, each object functions as an independent entity, capable of creating complex scenes through interaction with other objects. Our primary objective is to generate high-quality 3D objects, and while the current limitation exists, it does not diminish the overall quality and utility of our approach.
5. Conclusion